OpenStudy (anonymous):

Solve using the multiplication principle. 12c = -25

OpenStudy (anonymous):

Multiply both sides by the reciprocal of the coefficient of the variable. 12 is the coefficient; it's reciprocal is 1/12. So (1/12)12c = (1/12)(-25); c = -25/12.